OpenStudy (anonymous):

This is an example of an arithmetic series: 6 + 8 + 10 + 12 + . . . + (4 + 2n) + . . . True False

OpenStudy (btaylor):

true, because there is always a constant difference between the terms. ex: 8-6 = 10-8 = 12-10 = 2

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

Use : \[\large a_n = a + (n-1)d\]

OpenStudy (anonymous):

Here a = 6 and d = 8 - 6 = 2 \[a_n = 6 + (n-1)2 \implies 4 + 2n\] It is an Arithmetic Progression..